# DDIL: Improved Diffusion Distillation with Imitating Learning

### v1

@@ Abstract
Diffusion models excel at generative modeling (e.g., text-to-image) but sampling requires multiple denoising network passes, limiting practicality. Efforts such as progressive distillation or consistency distillation have shown promise by reducing the number of passes at the expense of quality of the generated samples. In this work we identify co-variate shift as one of reason for poor performance of multi-step distilled models from compounding error at inference time. To address co-variate shift, we formulate diffusion distillation within imitation learning (DDIL) framework and enhance training distribution for distilling diffusion models on both data distribution (forward diffusion) and student induced distributions (backward diffusion). Training on data distribution helps to diversify the generations by preserving marginal data distribution and training on student distribution addresses compounding error by correcting covariate shift. In addition, we adopt reflected diffusion formulation for distillation and demonstrate improved performance, stable training across different distillation methods. We show that DDIL consistency improves on baseline algorithms of progressive distillation (PD), Latent consistency models (LCM) and Distribution Matching Distillation (DMD2).
@@ 1 Introduction
Figure 1: Qualitative comparison of images generated with different distillation techniques.
Diffusion models, while capable of producing high-quality images, suffer from slow sampling times due to their iterative denoising process. To address this, distillation techniques have been proposed to reduce number of denoising steps. These techniques can be broadly categorized into trajectory-level (Luo et al., 2023; Meng et al., 2023; Salimans & Ho, 2021; Song et al., 2023) and distribution-matching approaches (Yin et al., 2023; 2024; Luo et al., 2024; Sauer et al., 2023) . While the former focuses on preserving the teacher’s trajectory at a per-sample level, the latter matches the marginal distribution.
Multi-step student models offer a promising approach in balancing quality and computational efficiency. However, they often face a critical challenge: covariate shift. This occurs when the distribution of noisy input latents that the student model encounters during training differs from that seen during inference by the student. This mismatch can significantly impact generation quality especially when the number of denoising steps are low. Recent works Kohler et al. (2024); Yin et al. (2024) only consider backward trajectories to obtain feedback on quality of generation, but these approaches are often agnostic to data distribution and can exhibit mode-collapse.
In this work, we identify ‘covariate shift’ as a critical factor that impacts the generation quality in multi-step distilled diffusion models. To address covariate shift and to preserve diversity, we introduce diffusion distillation within the imitation learning (DDIL) framework by improving the training distribution for distillation. We achieve this by incorporating both the data distribution (forward diffusion) and the student’s predictive distribution (backward trajectory at inference time). This approach combines the benefits of (1) Preserving Marginal Data Distribution: Training on the data distribution ensures the student model maintains the inherent statistical properties of the original data, and (2) Correcting Covariate Shift: Training on backward trajectories enables the student model to identify and adapt to covariate shifts, thereby improving the accuracy of score estimates, particularly in few-step settings. We illustrate instantiation of DDIL framework in context of progressive distillation in Figure 2. To this end, we make the following contributions:
@@
  • We propose a novel DDIL framework which enhances training distribution of the diffusion distillation within the dataset aggregation ‘DAgger’ framework by performing distillation on on both the data distribution (forward) and student induced distribution (backward trajectory), yielding improved aggregate predictive distribution and better coverage.
  • To enhance the stability of the distillation process in diffusion models, we adopt thresholding for both the teacher and student diffusion models to enforce the support of the data distribution with reflected diffusion Lou & Ermon (2023) for distillation. Consequently, this approach further mitigates covariate shift, leading to more substantial improvements when combined with DDIL
  • We demonstrate that our DDIL approach yields diverse samples and consistently improves on different distillation techniques like progressive distillation (PD), latent consistency distillation (LCM), distribution matching distillation (DMD2) in a computationally efficient framework.
@@ 2 Related Work
Diffusion distillation methods. Progressive distillation Salimans & Ho (2021); Meng et al. (2023) and many follow up works Li et al. (2023b); Berthelot et al. (2023) try to reduce the number of iterations of student model by forcing student to mimic multiple steps of the teacher. Consistency models Song et al. (2023); Luo et al. (2023); Ren et al. (2024) assume deterministic probabilistic flow at inference and enforce consistency in the data space for step-distillation. Additionally, recent work decomposes the diffusion trajectory into multiple segments like in progressive distillation and performs distillation within consistency formulation Kim et al. (2023). Instead of using the real data, methods such as BOOT Gu et al. (2023) consider bootstrapping in the student trajectory to generate samples of high quality and diversity. Liu et al. (2023) approximates the underlying map of the pretrained diffusion model as linear paths. While above trajectory level distillation techniques like progressive distillation and consistency-based approaches improve efficiency, the quality of the generated samples exhibits low visual fidelity.
Alternatively, diffusion distillation has been formulated in the distributional matching framework Yin et al. (2023); Luo et al. (2024); Yin et al. (2024); Salimans et al. (2024); Sauer et al. (2023; 2024). Within distribution matching approaches instead of matching teacher for each trajectory or particle like in previous class of methods, we try to match marginals of distilled student model and pretrained diffusion model. Further, adversarial loss has been applied to distillation approaches to improve the visual quality of the generated images Sauer et al. (2023; 2024); Lin et al. (2024). Most of distributional matching objectives like Sauer et al. (2024); Yin et al. (2023) are mode-seeking and looses on diversity. EM distillation Xie et al. (2024) addresses this by richer sampling with langevian MCMC to provide better target for distillation.
Reverse diffusion as Markov decision process. Policy gradient methods have recently gained traction in text-to-image generation with diffusion models by formulating the reverse diffusion process as a markov decision process (MDP) Fan et al. (2024); Xu et al. (2023a). Recent work Fan et al. (2024) proposes a policy gradient method for data distribution matching in diffusion models. Black et al. (2023) introduces a policy gradient algorithm with reward function that optimizes a diffusion model for downstream tasks. Yang et al. (2024) assumes a latent reward function of the reverse denoising process by emphasing the text and image alignment on the coarser steps of image generation. All these approaches have been applied to improve the alignment between the prompts and generated images for high-fidelity synthesis. In our work, we leverage the formulation of reverse process as MDP for step-distillation. This formulation allows interactively update the student model with the observations of the teacher model using dataset aggregation Ross et al. (2011).
@@ 3.1 Reverse Denoising Process as MDP
In imitation learning, an agent learns to perform tasks by observing and mimicking the behavior of the expert. An MDP in imitation learning models the next action based on the previous action and the current knowledge of the environment (Ke et al., 2021; Spencer et al., 2021). In general, an MDP is represented as $\langle\mathcal{S},\mathcal{A},P,\rho_{0}\rangle$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is the set of actions, $P(s^{\prime}|s,a)$ is a state transition kernel to transition from $s$ to $s^{\prime}$ under the action $a$ and $\rho_{0}$ is the set of initial states. An MDP produces a trajectory which is a sequence of state-action pairs $\tau=(s_{0},a_{0},s_{1},a_{1},...,a_{T},s_{T})$ over $T$ time steps.
We formalize the reverse process of the diffusion models as a finite horizon MDP (Black et al., 2023; Fan et al., 2024) with the policy $\pi_{\theta}$ (the diffusion model with parameters $\theta$) where the states and the actions are $s_{t}\coloneqq(\mathbf{x}_{t},t)$ and $a_{t}\coloneqq\mathbf{x}_{t-1}$ respectively. The transition dynamics is defined by $P(s_{t+1}|s_{t},a_{t})\coloneqq\delta(\mathbf{x}_{t-1},t)$ and $\rho_{0}(s)\coloneqq(\mathcal{N}(\mathbf{0},\mathbf{I}),T)$ denotes the initial state distribution. The trajectory $\tau$ becomes $(\mathbf{x}_{T},\mathbf{x}_{T-1},\ldots,\mathbf{x}_{0})$.
@@ 3.2 Co-variate Shift in Diffusion Models
With in iterative denoising steps of generation within backward trajectory of diffusion models, student’s current predictions determines what the student (learner) sees in next step within sequential setting, which is classic feedback loop [1] in imitation learning. So if student makes any mistake or has bad score estimate in one of early steps this discrepancy excarbates in later iterations and results in accumulation of error. This error results in change in input distribution (covariate shift) of latents between training time (forward diffusion) and latents student model encounters when it is unrolled in iterative fashion at generation i.e., backward trajectory. Exposure bias is another closely related line of work Li et al. (2023a) which also discusses change in input distribution w.r.t pretrained diffusion model and propose training-free methods to improve it. Our work primarily focuses on distilling diffusion models and how this shift effects distillation.
@@
Covariate shift is more prounced for distilled student diffusion model compared to pretrained diffusion model. To further clarify why covariate shift poses more of a challenge for the student model compared to the teacher model, we can consider inference as ancestral sampling (or annealing in score estimation). During generation i.e., within intermediate time-steps of backward trajectory of diffusion model, there is an implicit assumption that the marginal distributions between two consecutive denoising steps significantly overlap. While this is a reasonable assumption in continuous time diffusion models or when the number of denoising steps is sufficiently high for pretrained diffusion model, when considering a diffusion model with only few steps, this assumption does not hold. Consequently, any covariate shift would be more exacerbated for the student model, unlike the continuous time teacher mode.
@@
DAgger to mitigate Co-variate shift: Imitation learning has long been used to learn offline sequential tasks wherein a student model is trained from teachers’ demonstrations. Standard imitation learning also suffers from covariate shift in discrepancy in states visited by student and the teacher. Interactive methods such as DAgger (Ross et al., 2011) in Imitation learning augment training data by querying the teacher model on student’s states, thereby obtaining teacher’s feedback on student’s predictive distribution(backward trajecories). Building on the ideas of interactive methods in imitation learning, in our work we aim to improve training distribution for diffusion distillation.
@@
Imitation Learning as Distribution Matching: Notably, Ke et al. (2021) has shown that the imitation learning algorithms can be formalized as $f$-divergence minimization algorithms where the DAgger approach minimizes the total variation (TV) distance between the student and the teacher trajectory distributions. If $\rho_{\theta}(\tau)$ and $\rho_{\eta}(\tau)$ are teacher and student trajectory distributions respectively, then DAgger minimizes upperbound on total variation. DAgger achieves $O(T\epsilon)$ error compared to behavior cloning equivalent progressive distillation or any method with teacher forcing bounded by $O(T\epsilon^{2})$ error.
$D_{TV}(\rho_{\theta}(\tau),\rho_{\eta}(\tau))$ $\leq T\mathbb{E}_{s\sim\rho_{\eta}(s)}\left[D_{TV}(p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t),p_{\eta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)\right]$ (1)
@@ 3.3 Backward Trajectories for Distilling Diffusion Models
Backward trajectory distillation introduced in recent concurrent works like ImagineFlash(Kohler et al., 2024) and DMD2 (Yin et al., 2024), focuses solely on evaluating the quality of generated samples without considering the data distribution. Consequently, they lack a mechanism to prevent mode collapse and ensure diversity.
@@
While EM Distillation (Xie et al., 2024) addresses this by employing Langevin MCMC for a richer reverse process and mode-covering divergences, it still doesn’t explicitly incorporate the data distribution into its sampling prior during distillation.
@@
Progressive Distillation (Meng et al., 2023; Salimans & Ho, 2021)
ImagineFlash (Kohler et al., 2024)
@@ 4.1 Improving Training Distribution with DDIL
We introduce Diffusion Distillation with Imitation Learning (DDIL), a novel framework inspired by the DAgger algorithm from imitation learning to enhance the sampling distribution of intermediate noisy latents for distilling diffusion models. Diffusion model distillation involves two key considerations: (1) the training distribution of latent states encountered by the student model, and (2) the feedback mechanism employed during distillation. DDIL specifically focuses on improving the training distribution, remaining agnostic to the specific feedback mechanism utilized by different distillation techniques.
@@
To achieve this, DDIL strategically samples intermediate latent variables from three sources: (1) forward diffusion of the dataset, captured by the sampling prior $\beta_{frwd}$ (as illustrated in Algorithm 1); (2) backward trajectories (unrolled latents) from the student model, denoted by the sampling prior $\beta_{student\_bckwrd}$; and (3) backward trajectories from the teacher model, denoted by the sampling prior $\beta_{teacher\_bckwrd}$, which is particularly advantageous in data-free settings for preserving marginal data distribution. Combining these sampling strategies leads to improved distillation performance.
@@
DDIL is a unified training framework for distilling diffusion models w.r.t sampling prior for distillation. DDIL incorporates teacher feedback on student trajectories, aligning with the principles of DAgger Ross et al. (2011) in case of progressive distillation and latent consistency models (LCM). Furthermore, while methods like Kohler et al. (2024); Yin et al. (2024) perform distillation only on backward trajectories and don’t account for marginal data distribution during distillation, DDIL addresses this by consistently incorporating feedback from the chosen distillation algorithm on both forward and backward trajectories. Our flexible framework thus allows for improved training distribution to boost the performance of diffusion distillation methods.
@@
Algorithm 1 outlines a generalized framework for Diffusion Distillation with Imitation Learning (DDIL). This framework leverages a pre-trained diffusion model (teacher) and a student diffusion model, typically initialized with the teacher’s parameters. Additionally, access to real data is assumed, providing representative samples from the marginal data distribution during the distillation process. The framework necessitates specifying hyperparameters for both the teacher and student models, including their respective discretization schemes. For simplicity we assume DDIM solver in 1. Distillation proceeds by randomly selecting one of three methods for sampling intermediate noisy latent inputs to the student model. This selection is governed by user-defined sampling priors: $\beta_{frwd}$, $\beta_{teach\_bckwrd}$, and $\beta_{student\_bckwrd}$, which correspond to the three sources of intermediate latents previously discussed. The choice and updating of these sampling priors, denoted as $\beta_{i}$, can be tailored based on the training stage, objective function, and overall task goals.
@@
Let $q_{\eta}$ be the predictive distribution of distilled student diffusion model from its generated backward trajectories, then overall DDIL objective is
@@
Where $L_{\text{Distill}}$ can assume any objective based on chosen algorithm like progressive distillation, latent consistency distillation and distribution matching objective and $L_{\text{DDIL}}$ trains on both data distribution and backward trajectories.
@@
Reflected Diffusion Distillation: When distilling diffusion models either the teacher or student model might not satisfy implicit assumed support during distillation which could makes training unstable and require large batch sizes, etc. We adopt reflected diffusion models Lou & Ermon (2023) framework for distillation i.e., threshold teacher’s score estimate and/or student’s and improve stability of training, lower required batch size with improved performance.
Thresholding is applied to the teacher model’s estimates consistently across all investigated methods: progressive distillation, Latent Consistency Models (LCM), and DMD2. Furthermore, within the consistency distillation framework, the target derived from the student model is also threshold-ed. For DMD2, thresholding is applied to the score estimates of the pre-trained diffusion model, the fake critic, and the student model. Without thresholding our gradient feedback could be noisy and negatively impacting training stability.

### v2

# DDIL: Diversity Enhancing Diffusion Distillation with Imitation Learning

@@ Abstract
Distilling Diffusion models has enabled practical adoption of diffusion models by reducing the number of iterations but distillation techniques often could suffer from lack of diversity, quality, etc. In this work we enhance training distribution for distilling diffusion models by training on both data distribution (forward diffusion) and student induced distributions (reverse process at inference). We formulate diffusion distillation within imitation learning (DDIL) framework and identify co-variate shift i.e., difference in intermediate marginal distributions between training and inference leading to poor performance of multi-step distilled models from compounding error at inference time.
Training on data distribution helps to diversify the generations by preserving marginal data distribution and training on student distribution addresses compounding error by correcting covariate shift. In addition, we adopt reflected diffusion formulation for distillation and demonstrate improved performance, stable training across distillation methods resulting in 18.73 FID@30k on SSD1B with just batch size ‘7’ and 40K updates! We show that DDIL consistently improves on baseline algorithms of Distribution Matching Distillation (DMD2), consistency distillation (LCM) and progressive distillation (PD).
@@ 1 Introduction
Figure 1: DDIL consistently improves both sample quality (FID@30k) and diversity(Coverage) across distillation approaches. Performance shown for DMD2 applied to SSD1B, and Consistency/Progressive Distillation applied to SDv1.5. Coverage measures the extent to which the generated samples span the real data manifold.
Figure 2: Qualitative comparison of images generated with different distillation techniques.We can observe more coherent structure with DDIL compared to baselines DMD2 even with thresholding e.g., space station structure or motorcycle structure. All distilled models are trained on same dataset, batch size and evaluated on same seed and hence generations share characteristics
Diffusion models, while capable of producing high-quality images, suffer from slow sampling times due to their iterative denoising process. To address this, distillation techniques have been proposed to reduce number of denoising steps. These techniques can be broadly categorized into trajectory-level [22, 25, 32, 37] and distribution-matching approaches [43, 44, 23, 34] . While the former focuses on preserving the teacher’s trajectory at a per-sample level, the latter matches the marginal distribution.
Multi-step student models offer a promising approach in balancing quality and computational efficiency. However, multi-step student model also suffer from ‘covariate shift’, i.e., difference in real and student’s assumed marginal distributions at intermediate time-steps or noise-levels. This is largely due to training vs inference mismatch in diffusion models and the effect on performance become more pronounced in few-step regime of distilled models. Recent works [12, 44] address this issue by distilling on backward trajectories but ignore diversity of generations. There is significant room for improvement on diversity-preserving distillation and need for consistent reporting of relevant metrics and identify better metrics w.r.t diversity.
In this work, we formulate diffusion distillation within the imitation learning (DDIL) framework and address covariate shift and also identify implicit assumptions in training distributions across different distillation techniques potentially leading to loss of diversity of distilled models.We achieve this by incorporating both the data distribution (forward diffusion) and that student’s predictive distribution (backward trajectory at inference time).
This approach combines the benefits of (1) Preserving Marginal Data Distribution: Training on the data distribution ensures the student model maintains the inherent statistical properties of the original data, and (2) Correcting Covariate Shift: Training on backward trajectories enables the student model to identify and adapt to covariate shifts, thereby improving the accuracy of score estimates, particularly in few-step settings. We illustrate instantiation of DDIL framework in context of progressive distillation in Figure 3. To this end, we make the following contributions:
@@
  • We propose a novel DDIL framework which enhances training distribution of the diffusion distillation within the dataset aggregation ‘DAgger’ framework by performing distillation on both the data distribution (forward) and student induced distribution (backward trajectory at inference time), yielding improved aggregate predictive distribution.
  • To enhance the stability of the distillation process in diffusion models, we adopt thresholding for both the teacher and student diffusion models to enforce the support of the data distribution with reflected diffusion [20] for distillation. Consequently, this approach further mitigates covariate shift, leading to more substantial improvements when combined with DDIL
  • We demonstrate that DDIL can be integrated with various distillation approached, and DDIL consistently improves on both quality and diversity within computationally efficient framework. DDIL only requires maintaining a prioritized replay buffer and hence does not incur significant computational or memory overhead.
@@ 2 Related Work
Diffusion distillation methods. Progressive distillation [32, 25] and many follow up works [15, 1] try to reduce the number of iterations of student model by forcing student to mimic multiple steps of the teacher. Consistency models [37, 22, 28] assume deterministic probabilistic flow at inference and enforce consistency in the data space for step-distillation. Additionally, recent work decomposes the diffusion trajectory into multiple segments like in progressive distillation and performs distillation within consistency formulation [11]. Instead of using the real data, methods such as BOOT [5] consider bootstrapping in the student trajectory to generate samples of high quality and diversity. [18] approximates the underlying map of the pretrained diffusion model as linear paths. While above trajectory level distillation techniques like progressive distillation and consistency-based approaches improve efficiency, the quality of the generated samples exhibits low visual fidelity.
Alternatively, diffusion distillation has been formulated in the distributional matching framework [43, 23, 44, 33, 34, 35]. Within distribution matching approaches instead of matching teacher for each trajectory or particle like in previous class of methods, we try to match marginals of distilled student model and pretrained diffusion model. Further, adversarial loss has been applied to distillation approaches to improve the visual quality of the generated images [34, 35, 16]. Most of distributional matching objectives like [35, 43] are mode-seeking and looses on diversity. EM distillation [39] addresses this by richer sampling with langevian MCMC to provide better target for distillation.
Reverse diffusion as Markov decision process. Policy gradient methods have recently gained traction in text-to-image generation with diffusion models by formulating the reverse diffusion process as a markov decision process (MDP) [4, 40]. Recent work [4] proposes a policy gradient method for data distribution matching in diffusion models. [2] introduces a policy gradient algorithm with reward function that optimizes a diffusion model for downstream tasks. [42] assumes a latent reward function of the reverse denoising process by emphasing the text and image alignment on the coarser steps of image generation. All these approaches have been applied to improve the alignment between the prompts and generated images for high-fidelity synthesis. In our work, we leverage the formulation of reverse process as MDP for step-distillation. This formulation allows interactively update the student model with the observations of the teacher model using dataset aggregation [30].
@@ 3.1 Reverse Denoising Process as MDP
In imitation learning, an agent learns to perform tasks by observing and mimicking the behavior of the expert. An MDP in imitation learning models the next action based on the previous action and the current knowledge of the environment [10, 38]. In general, an MDP is represented as $\langle\mathcal{S},\mathcal{A},P,\rho_{0}\rangle$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is the set of actions, $P(s^{\prime}|s,a)$ is a state transition kernel to transition from $s$ to $s^{\prime}$ under the action $a$ and $\rho_{0}$ is the set of initial states. An MDP produces a trajectory which is a sequence of state-action pairs $\tau=(s_{0},a_{0},s_{1},a_{1},...,a_{T},s_{T})$ over $T$ time steps.
We formalize the reverse process of the diffusion models as a finite horizon MDP [2, 4] with the policy $\pi_{\theta}$ (the diffusion model with parameters $\theta$) where the states and the actions are $s_{t}\coloneqq(\mathbf{x}_{t},t)$ and $a_{t}\coloneqq\mathbf{x}_{t-1}$ respectively. The transition dynamics is defined by $P(s_{t+1}|s_{t},a_{t})\coloneqq\delta(\mathbf{x}_{t-1},t)$ and $\rho_{0}(s)\coloneqq(\mathcal{N}(\mathbf{0},\mathbf{I}),T)$ denotes the initial state distribution. The trajectory $\tau$ becomes $(\mathbf{x}_{T},\mathbf{x}_{T-1},\ldots,\mathbf{x}_{0})$.
@@ 3.2 Co-variate Shift in Diffusion Models
With in iterative denoising steps of generation within backward trajectory of diffusion models, student’s current predictions determines what the student (learner) sees in next step within sequential setting, which is classic feedback loop [38] in imitation learning. So if student makes any mistake or has bad score estimate in one of early steps this discrepancy exacerbates in later iterations and results in accumulation of error. This error results in change in input distribution (covariate shift) of latents between training time (forward diffusion) and latents student model encounters when it is unrolled in iterative fashion at generation i.e., backward trajectory. Exposure bias is another closely related line of work [14] which also discusses change in input distribution w.r.t pretrained diffusion model and propose training-free methods to improve it. Our work primarily focuses on distilling diffusion models and how this shift effects distillation.
@@
Covariate shift is more pronounced for distilled student diffusion model compared to pretrained diffusion model. To further clarify why covariate shift poses more of a challenge for the student model compared to the teacher model, we can consider inference as ancestral sampling (or annealing in score estimation). During generation i.e., within intermediate time-steps of backward trajectory of diffusion model, there is an implicit assumption that the marginal distributions between two consecutive denoising steps significantly overlap, which enables diffusion models to self-correct towards high likelihood. While overlap in marginal distributions of consecutive time-steps is a reasonable assumption in continuous time diffusion models or when the number of denoising steps are sufficiently high, when considering a diffusion model with only few steps, this assumption does not hold and any we might be querying student model in out-of-domain at inference time causing performance degradation.
Consequently, any covariate shift would be more exacerbated for the student model, unlike the continuous time teacher model. We do empirical analysis of covariate shift in 5.2
@@
DAgger to mitigate Co-variate shift: Imitation learning has long been used to learn offline sequential tasks wherein a student model is trained from expert or teacher demonstrations. Standard imitation learning also suffers from covariate shift, i.e., discrepancy in states visited by student and the teacher. Interactive methods such as DAgger [30] in Imitation learning augment training data by querying the teacher model on student’s states, thereby obtaining teacher’s corrective feedback on states that student has not seen in demonstration data. Building on the ideas of interactive methods in imitation learning, in our work we aim to improve training distribution for diffusion distillation.
@@
Imitation Learning as Distribution Matching: Notably, [10] has shown that the imitation learning algorithms can be formalized as $f$-divergence minimization algorithms where the DAgger approach minimizes the total variation (TV) distance between the student and the teacher trajectory distributions. If $\rho_{\theta}(\tau)$ and $\rho_{\eta}(\tau)$ are teacher and student trajectory distributions respectively, then DAgger minimizes upper bound on total variation. DAgger achieves $O(T\epsilon)$ error compared to behavior cloning equivalent progressive distillation or any method with teacher forcing bounded by $O(T\epsilon^{2})$ error.
@@ 3.3 Backward Trajectories for Distilling Diffusion Models
Backward trajectory distillation introduced in recent concurrent works like ImagineFlash[12] and DMD2 [44], focuses solely on evaluating the quality of generated samples without considering the data distribution. Consequently, they lack a mechanism to prevent mode collapse and ensure diversity.
@@
While EM Distillation [39] addresses this by employing Langevin MCMC for a richer reverse process and mode-covering divergences, it still doesn’t explicitly incorporate the data distribution into its sampling prior during distillation. Depending on right choice of divergence measure, we can consider better mode-preserving distillation objectives and can be integrated with DDIL.
@@
Progressive Distillation [25, 32]
ImagineFlash [12]
@@ 4.1 Improving Training Distribution with DDIL
We introduce Diffusion Distillation with Imitation Learning (DDIL), a novel framework inspired by the DAgger algorithm from imitation learning to enhance the sampling distribution of intermediate noisy latents for distilling diffusion models. Diffusion distillation involves two key considerations: (1) the training distribution of latent states encountered by the student model, and (2) the feedback mechanism employed during distillation. DDIL specifically focuses on improving the training distribution, remaining agnostic to the specific feedback mechanism utilized by different distillation techniques.
@@
To achieve this, DDIL strategically samples intermediate latent variables from three sources: (1) forward diffusion of the dataset, captured by the sampling prior $\beta_{frwd}$ (as illustrated in Algorithm 1); (2) backward trajectories (unrolled latents) from the student model, denoted by the sampling prior $\beta_{student\_bckwrd}$; and (3) backward trajectories from the teacher model, denoted by the sampling prior $\beta_{teacher\_bckwrd}$, which is particularly advantageous in data-free settings as a proxy in preserving marginal data distribution. Combining these sampling strategies leads to improved distillation performance.
DDIL is a unified training framework for distilling diffusion models w.r.t sampling prior for distillation. DDIL incorporates corrective feedback using $\beta_{student\_bckwrd}$ by preventing accumulation of error on student induced $z_{t}$ w.r.t distilled model. Specifically DDIL modifies progressive distillation (PD) and latent consistency models (LCM) by unrolling student and obtain (teacher’s) corrective feedback on student’s backward trajectories i.e., if a student observes less encountered $z_{t}^{s}$ then obtaining a feedback on this rare $z_{t}^{s}$ improves student’s estimate at $z_{t}^{s}$ but in teacher forced distillation methods like PD we only $z_{t}^{s}$ at inference and never at training.
@@
Furthermore, while methods like [12, 44] perform distillation on backward trajectories to mitigate covariate shift, they don’t account for marginal data distribution during distillation and hence could loose diversity. DDIL addresses this by consistently incorporating feedback from the chosen distillation algorithm on both forward and backward trajectories, i.e., $\beta_{frwd}$ or $\beta_{teacher\_bckwrd}$ are necessary during distillation to better be aware of data distribution which is not the case in methods like GANs,[44], etc. Our flexible framework thus allows for improved training distribution to boost the performance of diffusion distillation methods.
@@
Algorithm 1 outlines a generalized framework for Diffusion Distillation with Imitation Learning (DDIL). This framework leverages a pre-trained diffusion model (teacher) and a student diffusion model, typically initialized with the teacher’s parameters. Additionally, access to real data or equivalent prompts are assumed, providing representative samples from the marginal data distribution during the distillation process. The framework necessitates specifying hyper-parameters for both the teacher and student models, including their respective discretization schemes. For simplicity we assume DDIM solver in 1. Distillation proceeds by randomly selecting one of three methods for sampling intermediate noisy latent ‘inputs’ to the student model. This selection is governed by user-defined sampling priors: $\beta_{frwd}$, $\beta_{teach\_bckwrd}$, and $\beta_{student\_bckwrd}$, which correspond to the three sources of intermediate latents previously discussed. The choice and updating of these sampling priors, denoted as $\beta_{i}$, can be tailored based on the training stage, objective function, and overall task goals. If distillation is performed without image data i.e., data-free settings, then $\beta_{frwd}=0$ and $\beta_{teach\_bckwrd}$ acts as a proxy to sample from data distribution.
@@
Let $q_{\eta}(x)$ be aggregate predictive distribution of distilled student model from its generated trajectories. DDIL objective is to train on sampled latents from student’s predictive distribution $q_{\eta}(x)$ and data distribution $p_{data}(x)$
@@
Where $L_{\text{Distill}}$ can assume any objective based on chosen algorithm like progressive distillation, latent consistency distillation and distribution matching objective. This makes student model to match it’s $q_{\eta}(x)$ to $p_{data}(x)$ not just at $t=0$ but also at other intermediate time-steps and corresponding noisy marginal distributions/
@@
Reflected Diffusion Distillation: When distilling diffusion models either the teacher or student model might not satisfy implicit assumed support during distillation which could makes training unstable and require large batch sizes, etc. We adopt reflected diffusion models [20] framework for distillation i.e., threshold score estimates of teacher model and/or student model.
Static thresholding [31] is applied to the teacher model’s estimates consistently across all investigated methods: progressive distillation, Latent Consistency Models (LCM), and DMD2. Furthermore, within the consistency distillation framework, thresholding is also applied to the target derived from the student model. In case of DMD2, thresholding is applied to the score estimates of the pre-trained diffusion model, the fake critic, and the student model. Without thresholding, our gradient feedback could be noisy and negatively impacting training stability. We observe that when training DMD2 with small batch (7), without thresholding performance deteriorates and when we adopt thresholding can observe monotonic improvement in performance.